Solly Corporation produces a product for national distribution. Standards for the product are: • Materials: 12 ounces per unio $4,000 U o $4,000E o $4,125 U o $4,125F
Answer #1

Labor efficiency variance

= (SH - AH) * SR

= (1000*2 - 2500) * 8

= (2000 - 2500) * 8

= 4000 U

OPTION A is the answer